OpenStudy (anonymous):

During a hike, Vijay saw a total of 19 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 52 wings. Use a model to determine how many birds he saw. 9 10 11 12

OpenStudy (stamp):

During a hike, Vijay saw a total of 19 birds and butterflies. a total of 19 birds and butterflies. 19 birds and butterflies. 19 birds + butterflies. 19 = birds + butterflies birds have 2 wings and butterflies have 4 wings, he counted 52 wings

OpenStudy (stamp):

\[2\ \frac{wings}{bird}*birds=wings \ from\ birds\]\[4\frac{wings}{butterfly}*butterflies=butterfly\ wings\]or 2birds = bird wings 4butterflies = butterfly wings Add up the wings together. remember how he counted 52 wings. 2birds + 4 butterflies = 52 birds + butterflies = 19 Now solve for birds and butterflies.

OpenStudy (stamp):

I should say , to avoid confusion, 2*birds + 4*butterflies = 52 birds + butterflies = 19

OpenStudy (stamp):

Hint: to solve, use substitution.

OpenStudy (stamp):

birds = 19 - butterflies 2(19-butterflies) + 4butterflies = 52 38 - 2 butterflies + 4butterflies = 52 2 butterflies = 14 butterflies = 7 birds + 7 = 19 birds = 12